A superman is sitting on a tree 98m right with a baby he has rescued from the claws of a tiger. Unfortunately the child slips and fall with an initial velocity of 0ms^–². Superman realized what was happening 2 seconds later and flow to catch the child at a speed of 39.2ms^–¹. Calculate







(1) the time superman will fly before catching the child.







(2) the distance superman will fly before catching the child.

1
Expert's answer
2022-01-24T11:09:37-0500

The equation of motion of a superman

"y_s=39.2(t-2)"

The equation of motion of a child

"y_c=9.8t^2\/2"

At the instant when superman catches the child

"y_s=y_c"

"39.2(t-2)=9.8t^2\/2"

"t^2-8t+16=0"

(1) the time superman will fly before catching the child:

"\Delta t=4-2=2\:\rm s"

(2) the distance superman will fly before catching the child:

"y_s=39.2*2=78.4\:\rm m"